# Signed Networks: theory, methods, and applications

### v1

@@ 6.1 Mining signed networks from raw data
Figure 17: Data sources for signed network inference. The figure illustrates different types of data that can be used to infer signed networks, including bipartite projections, aggregated unsigned interactions, online textual or voting interaction data and co-edit patterns in Wikipedia.
@@ 6.1.1 Ranking data
A popular approach to infer signed relations relies on ego network rankings, where each participant is asked to rank others in terms of their preference to interact with them. One of the first well-known examples is the Newcomb dataset [newcombAcquaintanceProcess1961], in which a group of students was asked to rank the remaining group members across several time points. From such rankings, signed relations can be inferred by selecting thresholds that determine which positions in the ranking reflect positive, negative, or neutral links [doreianBriefHistoryBalance1997]. For instance, the top-ranked individuals might be interpreted as positive links, the bottom-ranked as negative, and the rest as neutral. The main limitation of this method is that details on the mapping from ordinal ranks to signed links (e.g. the choice of thresholds) can affect the resulting network structure.
However, as Hallinan [hallinanAnalysisIntransitivitySociometric1975] points out, fixing the number of allowed alters (e.g., limiting participants to only select a fixed number of positive and negative links or forcing the selection a posteriori from ranking data) can lead to artefacts in structural analysis. In particular, constraints on the number of declared links reduce the number of possible closed triangles, potentially underestimating the level of balance in the network. One proposed solution is to allow participants to list as many positive or negative links as they wish, which results in a more accurate reflection of the social structure and increases the chance of observing structurally balanced triads.
@@ 6.1.2 Unsigned bipartite networks
Many real-world datasets are naturally bipartite [^29] The terms unipartite and one-mode (respectively, bipartite and two-mode) are often used interchangeably in the literature., meaning they contain two types of entities (called agents and artefacts) where connections exist only across the two sets. Examples include authors writing papers, legislators sponsoring bills, or customers reviewing products they bought. From these data, researchers frequently infer a co-occurrence, unipartite network, where two agents are connected if they share artefacts (e.g., two senators co-sponsoring the same bill). This process, known as bipartite projection, typically results in a positively weighted network, with edge weights reflecting the number of shared artefacts. Mathematically, given a bipartite network represented as the rectangular matrix $B\in\{0,1\}^{n_{1}\times n_{2}}$, where each row represents an agent and each column represents an artefact, the one-mode projection onto the set of agents is defined as $W=BB^{\top},$ where $W_{ij}$ counts the number of shared artifacts between agents $i$ and $j$.
@@
This raw co-occurrence count method has several problems, as the projected network is often very dense, and it usually contains spurious data that does not really reflect the relation between agents. For example, if two senators co-sponsor a bill that everyone supports, that connection does not provide meaningful information about their relationship. Moreover, if the original bipartite network is unsigned, its projection will only include positive edges, while a signed representation could sometimes better capture the underlying relationships. Therefore, it is important to implement a statistical framework to test which connections are stronger or weaker than expected by chance. This process is called backbone extraction. Here, we outline the most popular approaches; for more details, please refer to [nealBackboneBipartiteProjections2014].
@@
The simplest way to extract a backbone is to apply a global threshold, removing all connections with weights below the defined value. This method assumes that raw co-occurrence counts are directly comparable across all pairs of agents. Its main advantage is its easy implementation, no need for statistical modelling, and its computational efficiency. However, it suffers from important limitations  [nealBackboneBipartiteProjections2014], such as the arbitrary choice of the threshold, as different thresholds can yield very different network structures. A low threshold produces a dense, almost complete graph, while a high threshold may leave very few edges. Moreover, the resulting network is often biased. For example, thresholded graphs typically have high clustering, even when the original bipartite network is entirely random [zalesky2012use], making properties such as heavy-tailed distributions likely to be statistical artefacts.
@@
More principled approaches compare the observed co-occurrences to what would be expected under a random null model (see Section 4.1). The null models used for backbone extraction often preserve the degree of each agent and each artefact. The constraints can be enforced either exactly, such as the Fixed Degree Sequence Model (FDSM) [zweigSystematicApproachOnemode2011], or drawn from a degree distribution, such as the Stochastic Degree Sequence Model (SDSM) [nealBackboneBipartiteProjections2014]. The former model has been used, for instance, to infer signed social networks from the bipartite projection of play interactions among preschoolers [nealInferringSignedNetworks2022], while the latter has been applied to extract signed networks of political relationships among members of the U.S. Congress [nealSignTimesWeak2020].
@@
In summary, both the FDSM and the SDSM test whether an observed co-occurrence between two agents is larger or smaller than expected under a degree-preserving (or degree-distribution-preserving) null model. The SDSM approach is typically the preferred one, since it is faster, scales better, and avoids over-conditioning the null model (see  [nealComparingAlternativesFixed2021] for a more in-depth comparison of the different models). A ready-to-use implementation of this method with different null model choices can be found in the R backbone package [nealBackbonePackageExtract2022]. A modified version of SDSM that incorporates edge constraints is presented in [nealStochasticDegreeSequence2024], and a recent extension by Gallo et al. [gallo2025statistically] adapts the method to bipartite signed networks.
Another relevant line of work has explored the inference of signed links from a bipartite structure of behavioural data. Yang et al. [yangFriendFrenemyPredicting2012] propose the Behaviour–Relation Interplay (BRI) model, which predicts the sign of social links, in this case trust or distrust, by jointly modelling user behaviour (agents) with artefacts or items (e.g., news content) and social interactions. The framework combines a series of latent factor models for user–item decisions and social relations to determine edge signs automatically, even in unsupervised settings. The fundamental assumption is that behavioural agreement is an indicator of positive links and disagreement is a signal of negative ones. Based on this assumption, the model captures the mutual influence between behaviour and relationships. They test their method on Epinion and Yahoo! and show that it provides stronger predictive power for behavioural modelling than standard homophily-based approaches.
@@ 6.1.3 Aggregates of unsigned interaction data
An alternative approach that infers weighted signed relations is the one from [andresReconstructingSignedRelations2023]. This method takes as input a collection of unsigned and repeated interactions between individuals instead of a bipartite structure. Their method relies on the assumption that, in social systems, a statistical over-representation of unsigned interactions signals a positive relation, while an under-representation signals a negative relation. It thus requires data on interactions and partial data on ground truth relations. Similar to the SDSM [nealBackboneBipartiteProjections2014], this method works by comparing the observed interaction counts with a network null model, in this case, the hypergeometric ensemble of random graphs (HypE) [casiraghiConfigurationModelsUrn2021]. HypE incorporates the heterogeneity in the activities and attractiveness of individuals, and has the advantage of being analytically tractable.
@@
Based on this null model, a weighted signed edge between two individuals is assigned according to the marginal probability that the observed number of interactions is higher or lower than its expected value. To determine the signed relation, these probabilities are weighted with community-dependent variables $a$ and $b$, learned through the declared ground truth relations in the datasets. This method is applied to five datasets with different features (size, directed/undirected, type of interactions, form of surveyed relations, etc.): a windsurfer community [freemanHumanSocialIntelligence1988], the known Zachary’s Karate Club dataset [zacharyInformationFlowModel1977], face-to-face interactions in a French high school [mastrandreaContactPatternsHigh2015], the Nethealth project [liuNetworkAnalysisNetHealth2018], and Epinions [hamedaniTrustRecEffectiveApproach2021].
@@
Similarly, in [gursoyExtractingSignedBackbone2021], they provide a method to extract the signed backbone of intrinsically dense weighted networks, which assumes that all nodes are aware of other nodes [^30] Because of this assumption, this method might be inappropriate for tasks like extracting signed networks from large-scale social media data.. They show applications to datasets related to migration, voting, temporal interaction and species similarity networks. The method extracts positive and negative links whose weights significantly and substantially deviate from the random expectation under a null model that preserves node strength. They introduce two filters to identify backbone links: the significance filter, which removes links not significantly different from expectation (using separate thresholds for positive and negative cases); and the vigour filter, which applies stricter thresholds based only on link intensity. This method accepts both directed and undirected networks. Their software is available in Python and R as signed.backbones.
@@ 6.1.4 Sparse online interactions
Most of the methods explained above are centred on extracting signed networks of personal relations in systems where it is assumed that all individuals are aware of the existence of other individuals, and then choose to interact more or less with each of them depending on the nature of their relation. In these settings, the absence of an edge between two nodes usually implies a neutral relationship between the two. This assumption correctly applies to elite political systems, international relations or co-authorship networks of small fields. However, it becomes problematic when applied to relations formed on social media. This is because in large-scale digital spaces, it is difficult to distinguish if a non-relation is the result of a conscious decision to avoid the other user or simply of a lack of exposure due to the algorithm or the user activity. This generates two types of neutral relations: those in which there has been interaction, but it does not lean towards positive or negative, and the cases in which there has been no opportunity for the interaction to happen. Yet both of these situations are categorised as non-edge. Moreover, in these platforms, the decision of what content or users one is exposed to is tightly defined by recommendation algorithms, which may have a strong effect on the final interactions and how these are distributed along the network. Finally, these interactions are also platform-specific: in some cases, we can find explicitly signed data, text reply data, or endorsement only, such as likes, retweets, or follows. The choice of how to translate these data structures to signed networks is then particular to each study, and we can find a large variation in collection and inference methods.
@@
One such method uses a Bayesian model that treats each signed interaction as a Bernoulli random variable [fraxanetUnpackingPolarizationAntagonism2024]. This model incorporates a beta prior to account for platform-level asymmetries between positive and negative interactions. Similar to the significance and vigour filter in [gursoyExtractingSignedBackbone2021], given the posterior probability, only pairs with a clear bias towards one type of interaction and low variance will form a signed relationship, which will be determined by the direction of the bias.
@@
A common approach to generate signed data from social platform interactions is to extract the agreement and disagreement from text, such as interactions through comments, with (Large) Language Models. In [pougue-biyongDEBAGREEMENTCommentreplyDataset], they introduce an annotated comment-reply dataset from Reddit, with labels of agree, disagree, and neutral, finding that state-of-the-art pre-trained models struggle most at capturing neutral interactions. Interestingly, they also propose as future work the combination of unsigned network data and text annotations to perform edge sign predictions with Graph Neural Networks (see Section 4.3).
Another work [keucheniusWhyItImportant2021] uses the fastText algorithm for text classification to classify Twitter mentions into positive and negative, and uses a majority vote system to infer relations between users. They show that, compared to analysing the Twitter debate with just unsigned interactions, the addition of antagonistic interactions enriches the understanding of the online debate by allowing for the separation of conflict and indifference.
Also using Twitter data, [tacchiSignedEgoNetwork2022] infer signed interactions from sentiment analysis using the VADER framework [huttoVADERParsimoniousRuleBased2014]. Their approach infers relations relying on the psychological theory of the golden interaction threshold, in which one relationship is considered positive as long as there is one or fewer negative interactions per five positive interactions (i.e. roughly 17%).
@@
Wikipedia has been the source of numerous studies that extract signed networks from co-editing behaviour and user interactions [brandesNetworkAnalysisCollaboration2009, sepehriradLeveragingEditorCollaboration2012, sepehri-radIdentifyingControversialWikipedia2015, nakamuraClusteringEditorsWikipedia2013, maniusilviuBuildingSignedNetwork, bogdanovCommunityDiscoverySigned2010, lernerFreeEncyclopediaThat2020]. Some of these approaches focus on building signed networks based on text interactions and try to obtain a clustering of the contributors to analyse their relationships [brandesNetworkAnalysisCollaboration2009, nakamuraClusteringEditorsWikipedia2013]. Others incorporate both textual and behavioural signals, including interactions across articles and adminship elections, to construct signed and directed networks used for tasks such as link prediction or article quality estimation [maniusilviuBuildingSignedNetwork].
Furthermore, Sepehri-Rad et al. [sepehriradLeveragingEditorCollaboration2012, sepehri-radIdentifyingControversialWikipedia2015] used co-editing patterns to predict user attitudes and detect controversial material within articles. They define individual, directional, and mutual features based on collaborators who co-edit the same articles, and use the adminship votes as ground truth labels to train a model that detects which features are more predictive of the “real” relationship amongst contributors. Other less conventional approaches assess the sign of the interaction based on the direction of the change in distance within a latent space of a topic model (i.e. if user $A$ moved the paragraph in the same direction as user $B$, they agree) [bogdanovCommunityDiscoverySigned2010].
Lastly, [lernerFreeEncyclopediaThat2020] introduced signed networks based on undo and re-do actions between users and explored the effects of balance and reputation theories on editor interactions. The authors also considered potential actions to account for implicit interactions (e.g. a text from A could have been deleted by B, but it has not) with a decay weight. The relation between two users was then explained by latent probabilities specified by logistic regression models based on past interactions.
@@ 6.2 Common data sources
In the following, we describe publicly available datasets of signed networks. We focus on four categories: in-person social communities, online platforms, political systems and country systems. Other relevant research fields leveraging signed networks, e.g. ecology, have been discarded due to the unavailability of public databases in signed network format. Summarised information can be found in Table A5 of the Appendix, alongside links to online data repositories.

### v2

@@ 6.1 Mining signed networks from raw data
Figure 17: Data sources for signed network inference. The figure illustrates different types of data that can be used to infer signed networks, including bipartite projections, aggregated unsigned interactions, online textual or voting interaction data, and co-edit patterns in Wikipedia.
@@ 6.1.1 Ranking data
A popular approach to infer signed relations relies on ego network rankings, where each participant is asked to rank others in terms of their preference to interact with them. One of the first well-known examples is the Newcomb dataset [292], in which a group of students was asked to rank the remaining group members across several time points. From such rankings, signed relations can be inferred by selecting thresholds that determine which positions in the ranking reflect positive, negative, or neutral links [94]. For instance, the top-ranked individuals might be interpreted as positive links, the bottom-ranked as negative, and the rest as neutral. The main limitation of this method is that details on the mapping from ordinal ranks to signed links (e.g., the choice of thresholds) can affect the resulting network structure.
However, as Hallinan [161] points out, fixing the number of allowed alters (e.g., limiting participants to only select a fixed number of positive and negative links or forcing the selection a posteriori from ranking data) can lead to artefacts in structural analysis. In particular, constraints on the number of declared links reduce the number of possible closed triangles, potentially underestimating the level of balance in the network. One proposed solution is to allow participants to list as many positive or negative links as they wish, which results in a more accurate reflection of the social structure and increases the chance of observing structurally balanced triads.
@@ 6.1.2 Unsigned bipartite networks
Many real-world datasets are naturally bipartite [^29] The terms unipartite and one-mode (respectively, bipartite and two-mode) are often used interchangeably in the literature., meaning they contain two types of entities (called agents and artefacts) where connections exist only across the two sets. Examples include authors writing papers, legislators sponsoring bills, or customers reviewing products they bought. From these data, researchers frequently infer a co-occurrence, unipartite network, where two agents are connected if they share artefacts (e.g., two senators co-sponsoring the same bill). This process, known as bipartite projection, typically results in a positively weighted network, with edge weights reflecting the number of shared artefacts. Mathematically, given a bipartite network represented as the rectangular matrix $B\in\{0,1\}^{n_{1}\times n_{2}}$, where each row represents an agent, and each column represents an artefact, the one-mode projection onto the set of agents is defined as $W=BB^{\top},$ where $W_{ij}$ counts the number of shared artifacts between agents $i$ and $j$.
@@
This raw co-occurrence count method has several problems, as the projected network is often very dense, and it usually contains spurious data that does not really reflect the relation between agents. For example, if two senators co-sponsor a bill that everyone supports, that connection does not provide meaningful information about their relationship. Moreover, if the original bipartite network is unsigned, its projection will only include positive edges, while a signed representation could sometimes better capture the underlying relationships. Therefore, it is important to implement a statistical framework to test which connections are stronger or weaker than expected by chance. This process is called backbone extraction. Here, we outline the most popular approaches; for more details, please refer to [291].
@@
The simplest way to extract a backbone is to apply a global threshold, removing all connections with weights below the defined value. This method assumes that raw co-occurrence counts are directly comparable across all pairs of agents. Its main advantage is its easy implementation, no need for statistical modelling, and its computational efficiency. However, it suffers from important limitations  [291], such as the arbitrary choice of the threshold, as different thresholds can yield very different network structures. A low threshold produces a dense, almost complete graph, while a high threshold may leave very few edges. Moreover, the resulting network is often biased. For example, thresholded graphs typically have high clustering, even when the original bipartite network is entirely random [411], making properties such as heavy-tailed distributions likely to be statistical artefacts.
@@
More principled approaches compare the observed co-occurrences to what would be expected under a random null model (see Section 4.1). The null models used for backbone extraction often preserve the degree of each agent and each artefact. The constraints can be enforced either exactly, such as the Fixed Degree Sequence Model (FDSM) [421], or drawn from a degree distribution, such as the Stochastic Degree Sequence Model (SDSM) [291]. The former model has been used, for instance, to infer signed social networks from the bipartite projection of play interactions among preschoolers [286], while the latter has been applied to extract signed networks of political relationships among members of the U.S. Congress [289].
@@
In summary, both the FDSM and the SDSM test whether an observed co-occurrence between two agents is larger or smaller than expected under a degree-preserving (or degree-distribution-preserving) null model. The SDSM approach is typically the preferred one, since it is faster, scales better, and avoids over-conditioning the null model (see  [287] for a more in-depth comparison of the different models). A ready-to-use implementation of this method with different null model choices can be found in the R backbone package [290]. A modified version of SDSM that incorporates edge constraints is presented in [288], and a recent extension by Gallo et al. [137] adapts the method to bipartite signed networks.
Another relevant line of work has explored the inference of signed links from a bipartite structure of behavioural data. Yang et al. [405] propose the Behaviour–Relation Interplay (BRI) model, which predicts the sign of social links, in this case trust or distrust, by jointly modelling user behaviour (agents) with artefacts or items (e.g., news content) and social interactions. The framework combines a series of latent factor models for user–item decisions and social relations to determine edge signs automatically, even in unsupervised settings. The fundamental assumption is that behavioural agreement is an indicator of positive links and disagreement is a signal of negative ones. Based on this assumption, the model captures the mutual influence between behaviour and relationships. They test their method on Epinion and Yahoo! and show that it provides stronger predictive power for behavioural modelling than standard homophily-based approaches.
@@ 6.1.3 Aggregates of unsigned interaction data
A particularly challenging case of signed network reconstruction arises when the observed data consist of contact frequencies rather than direct measures of social relationships. In this case, inference methods take as input a collection of unsigned and repeated interactions between individuals instead of a bipartite structure. They usually rely on the assumption that, in social systems, a statistical over-representation of unsigned interactions signals a positive relation, while an under-representation signals a negative relation. In such settings, the absence of interaction between two individuals is inherently ambiguous: it may arise from limited opportunity for contact rather than deliberate avoidance, and failing to account for this distinction risks misclassifying absent ties as negative ones. [120] address this issue with a Bayesian generative model that encodes interaction opportunities through latent interaction groups, thereby separating chance non-contact from socially meaningful avoidance. Their MCMC-based inference framework reconstructs signed networks while quantifying uncertainty in inferred edge signs, and performs especially well in detecting negative edges relative to baseline methods on synthetic benchmarks. Applied to face-to-face contact data from a French high school, the approach recovers signed structures consistent with questionnaire-derived friendship patterns, with posterior predictive checks supporting the adequacy of the model.
@@
An alternative approach that infers weighted signed relations is the one from [14]. Their method requires data on interactions and partial data on ground truth relations. Similar to the SDSM [291], this method works by comparing the observed interaction counts with a network null model, in this case, the hypergeometric ensemble of random graphs (HypE) [58]. HypE incorporates the heterogeneity in the activities and attractiveness of individuals, and has the advantage of being analytically tractable.
@@
Based on this null model, a weighted signed edge between two individuals is assigned according to the marginal probability that the observed number of interactions is higher or lower than its expected value. To determine the signed relation, these probabilities are weighted with community-dependent variables $a$ and $b$, learned through the declared ground truth relations in the datasets. This method is applied to five datasets with different features (size, directed/undirected, type of interactions, form of surveyed relations, etc.): a windsurfer community [133], the known Zachary’s Karate Club dataset [410], face-to-face interactions in a French high school [263], the Nethealth project [247], and Epinions [328].
@@
Similarly, in [159], they provide a method to extract the signed backbone of intrinsically dense weighted networks, which assumes that all nodes are aware of other nodes [^30] Because of this assumption, this method might be inappropriate for tasks like extracting signed networks from large-scale social media data.. They show applications to datasets related to migration, voting, temporal interaction, and species similarity networks. The method extracts positive and negative links whose weights significantly and substantially deviate from the random expectation under a null model that preserves node strength. They introduce two filters to identify backbone links: the significance filter, which removes links not significantly different from expectation (using separate thresholds for positive and negative cases); and the vigour filter, which applies stricter thresholds based only on link intensity. This method accepts both directed and undirected networks. Their software is available in Python and R as signed.backbones.
@@ 6.1.4 Sparse online interactions
Most of the methods explained above are centred on extracting signed networks of personal relations in systems where it is assumed that all individuals are aware of the existence of other individuals, and then choose to interact more or less with each of them depending on the nature of their relation. In these settings, the absence of an edge between two nodes usually implies a neutral relationship between the two. This assumption correctly applies to elite political systems, international relations, or co-authorship networks of small fields. However, it becomes problematic when applied to relations formed on social media. This is because in large-scale digital spaces, it is difficult to distinguish if a non-relation is the result of a conscious decision to avoid the other user or simply of a lack of exposure due to the algorithm or the user activity. This generates two types of neutral relations: those in which there has been interaction, but it does not lean towards positive or negative, and the cases in which there has been no opportunity for the interaction to happen. Yet both of these situations are categorised as non-edge. Moreover, in these platforms, the decision of what content or users one is exposed to is tightly defined by recommendation algorithms, which may have a strong effect on the final interactions and how these are distributed along the network. Finally, these interactions are also platform-specific: in some cases, we can find explicitly signed data, text reply data, or endorsement only, such as likes, retweets, or follows. The choice of how to translate these data structures to signed networks is then particular to each study, and we can find a large variation in collection and inference methods.
@@
One such method uses a Bayesian model that treats each signed interaction as a Bernoulli random variable [131]. This model incorporates a beta prior to account for platform-level asymmetries between positive and negative interactions. Similar to the significance and vigour filter in [159], given the posterior probability, only pairs with a clear bias towards one type of interaction and low variance will form a signed relationship, which will be determined by the direction of the bias.
@@
A common approach to generate signed data from social platform interactions is to extract the agreement and disagreement from text, such as interactions through comments, with (Large) Language Models. In [316], they introduce an annotated comment-reply dataset from Reddit, with labels of agree, disagree, and neutral, finding that state-of-the-art pre-trained models struggle most at capturing neutral interactions. Interestingly, they also propose as future work the combination of unsigned network data and text annotations to perform edge sign predictions with Graph Neural Networks (see Section 4.3).
Another work [209] uses the fastText algorithm for text classification to classify Twitter mentions into positive and negative, and uses a majority vote system to infer relations between users. They show that, compared to analysing the Twitter debate with just unsigned interactions, the addition of antagonistic interactions enriches the understanding of the online debate by allowing for the separation of conflict and indifference.
Also using Twitter data, [373] infer signed interactions from sentiment analysis using the VADER framework [196]. Their approach infers relations relying on the psychological theory of the golden interaction threshold, in which one relationship is considered positive as long as there is one or fewer negative interactions per five positive interactions (i.e., roughly 17%).
@@
Wikipedia has been the source of numerous studies that extract signed networks from co-editing behaviour and user interactions [50, 347, 348, 282, 256, 42, 229]. Some of these approaches focus on building signed networks based on text interactions and try to obtain a clustering of the contributors to analyse their relationships [50, 282]. Others incorporate both textual and behavioural signals, including interactions across articles and adminship elections, to construct signed and directed networks used for tasks such as link prediction or article quality estimation [256].
Furthermore, Sepehri-Rad et al. [347, 348] used co-editing patterns to predict user attitudes and detect controversial material within articles. They define individual, directional, and mutual features based on collaborators who co-edit the same articles, and use the adminship votes as ground truth labels to train a model that detects which features are more predictive of the “real” relationship amongst contributors. Other, less conventional approaches assess the sign of the interaction based on the direction of the change in distance within a latent space of a topic model (i.e., if user $A$ moved the paragraph in the same direction as user $B$, they agree) [42].
Lastly, [229] introduced signed networks based on undo and re-do actions between users and explored the effects of balance and reputation theories on editor interactions. The authors also considered potential actions to account for implicit interactions (e.g, a text from A could have been deleted by B, but it has not) with a decay weight. The relation between two users was then explained by latent probabilities specified by logistic regression models based on past interactions.
@@ 6.2 Common data sources
In the following, we describe publicly available datasets of signed networks. We focus on four categories: in-person social communities, online platforms, political systems, and country systems. Other relevant research fields leveraging signed networks, e.g. ecology, have been discarded due to the unavailability of public databases in signed network format. Summarised information can be found in Table A5 of the Appendix, alongside links to online data repositories.